A gas occupies a certain volume at 27 degree celcius. If it is heated at constant pressure,it volume is exactly doubled at a temperature of---
1
Expert's answer
2012-03-30T07:01:38-0400
Let's use the ideal gas law:

P*V = v*R*T.

Let assume T0 = 27° C. As pressure P is constant, then

2*P*V = v*R*T1 ==> T1 = 2*P*V/(v*R) = 2*(P*V)/(v*R) = 2*(v*R*T0)/(v*R) = 2*T0.

So, the volume doubles when the temperature doubles. Here doubled temperature is
T1 = 2*27 = 54° C.